Solve and check
a) log(x+2)+log(x-1)=1
b) log(3x+2)+log(x-1) = 2
c) 2log(x-1) = 2+log 100
Can you please help me out? thanks so much in advance:)
Can you show the work too because I will understand better thankyou:)

Answer by mananth(16946)

log(x+2)+log(x-1)=1

log((x+2)(x-1))=1
(x+2)(x-1)=10^1
x^2+x-2=10
x^2+x-12=0
(x+4)(x-3)=0
x=3 ignore negative


b) log(3x+2)+log(x-1) = 2
c) 2log(x-1) = 2+log 100